4/a + 9/b =-1
1/a + 6/b = 1

solve the system of equations

Answer by lwsshak3(11628)

4/a + 9/b =-1
1/a + 6/b = 1
**
4/a + 9/b =-1
1/a + 6/b = 1
multiply 2nd eq. by 4
4/a + 24/b=4
4/a + 9/b =-1
subtract
15/b=5
b=15/5=3
..
1/a + 6/b = 1
1/a + 6/3 = 1
1/a + 2 = 1
1/a=-1
a=1/-1=-1